OpenStudy (anonymous):

help. find the sixteenth term of the arithmetic sequence whose eleventh term is 96 and whose seventeenth term is 150

OpenStudy (anonymous):

Common difference is 9 since (150 - 96) / 6 "additions" from 11th to 17th = 54 / 6 = 9 If 11th term is 96 sequence must start at 6 since 96 - 9 * 10 is 6 (there are 10 terms before the 96) So then the equation is a(n) = 6 + (n-1)(9) So a(16) = 6 + 15(9) = 141 Or just subtract 9 from 150 to get 141 Here's the first 17 terms: 6 15 24 33 42 51 60 69 78 87 96 105 114 123 132 141 150

OpenStudy (anonymous):

thank you so much